# LLMind: Orchestrating AI and IoT with LLMs for Complex Task Execution

### v2

@@ Abstract
In this paper, we introduce LLMind, an AI framework that utilizes large language models (LLMs) as a central orchestrator. The framework integrates LLMs with domain-specific AI modules, enabling IoT devices to collaborate effectively in executing complex tasks. The LLM engages in natural conversations with human users via a user-friendly social media platform to come up with a plan to execute complex tasks. In particular, the execution of a complex task, which may involve the collaborations of multiple domain-specific AI modules and IoT devices, is realized through a control script. The LLM generates the control script using a Language-Code transformation approach based on finite-state machines (FSMs). The framework also incorporates semantic analysis and response optimization techniques to enhance speed and effectiveness. Ultimately, this framework is designed not only to innovate IoT device control and enrich user experiences but also to foster an intelligent and integrated IoT device ecosystem that evolves and becomes more sophisticated through continuing user and machine interactions.
I Introduction
@@
Large Language Model (LLM) has recently attracted tremendous attention from both academia and industry. It is an advanced generative artificial intelligence designed to comprehend and generate human-like text based on extensive training data. In addition to excellent language skills, LLM also exhibits a high level of reasoning and planning capability [1, 2]. Moreover, the remarkable feature of in-context learning substantially broadens the range of potential application scenarios for LLM, inspiring the design of diverse general-purpose AI agents.
@@
Our daily lives are inundated with an abundance of IoT devices, ranging from smart desk lamps to bulky smart home appliances such as refrigerators and washing machines. The emergence of LLM presents a transformative opportunity to revolutionize the development of IoT devices by empowering them with advanced intelligence. Imagine an LLM-based AI agent coordinating diverse IoT devices to collaboratively accomplish intricate tasks. LLM’s remarkable planning and scheduling flexibility surpasses scripted intelligence. Its in-context learning ability enables swift adaptation and optimal resource utilization for efficient task solutions. Consequently, users can effortlessly express their needs without concerns about operational details.
@@
Integrating LLMs into IoT device control presents a number of challenges despite its promising potential. First, the LLM cannot handle specialized tasks like object detection or facial recognition, while there are already AI models excelling in these domains. The first challenge lies in combining the power of fragmented AI modules with the LLM to handle various complex tasks. Furthermore, due to the LLM’s slow inference speed and high computational costs, it is crucial to investigate methods that can enhance the system’s response speed and overall efficiency. Another significant challenge involves the effective control of IoT devices by the AI agent. The LLM must accurately learn the diverse functionalities and operational characteristics of various IoT devices. Additionally, the design of efficient interaction mechanisms that enable users to naturally and intuitively express their needs to the LLM-based AI agent presents a further challenge.
In this work, we propose LLMind, a framework that transforms conventional IoT device control by integrating LLM as the orchestrator for accomplishing complex tasks. This integration leads to the construction of an LLM-empowered IoT ecosystem. Through novel approaches, our framework effectively overcomes the aforementioned challenges.
@@
First, the agent integrates domain-specific AI modules, with the LLM acting as the task allocation governor. This incorporation enables the general-purpose LLM to effectively manage specific tasks by invoking specialized AI modules, enhancing system flexibility and practicality. Moreover, by incorporating new AI modules, the AI agent acquires new capabilities and continues to evolve.
@@
Second, to enhance the response speed and overall efficiency of the agent, we utilize a semantic analysis NLP model to classify and locally store LLM-generated plans (control scripts) for specific user commands. By identifying similar instructions in historical records, even with varying tones or expressions, the agent can retrieve validated solutions, thus avoiding complete replanning. This approach enables the system to be more and more skillful through interactions.
@@
Third, the agent leverages the LLM for complex task planning and corresponding control script generation using a novel finite-state machine approach, enabling efficient and precise invocation of AI modules and IoT devices. This methodology significantly improves accuracy and success rates.
@@
We refer to the intelligence level attained by our framework as Artificial Complex Intelligence (ACI), which serves as a bridge between Artificial Narrow Intelligence (ANI) [3] and Artificial General Intelligence (AGI) [4], as illustrated in Fig. 1. ACI integrates the capabilities of various ANI modules and undergoes continuous evolution towards AGI. Our framework goes beyond mere device control, providing a comprehensive blueprint for a future intelligent, integrated, and collaborative IoT device ecosystem empowered by LLM.
II Framework Design
@@
This section presents the framework design. Our framework, depicted in Fig. 2, consists of five components: user interface, LLM, coordinator, AI modules, and IoT devices. The operational workflow of a demo system is illustrated in the flowchart presented in Fig. 3.
Figure 3: System operational process flowchart
Users can interact with the system by sending text messages through social media software. The LLM promptly responds to confirm message receipt. It is important to note that the system filters user messages. For casual chats and greetings, the system provides text replies without further processing. However, for specific requests and commands, the system formulates plans and strives to fulfill the user’s instructions using integrated AI modules and connected IoT devices.
After receiving a user instruction, the system prioritizes enhancing efficiency by searching for validated and feasible historical scripts stored locally in the coordinator’s long-term memory. This process utilizes NLP techniques to locate previously issued user commands with similar semantic meanings. If a matching control script is found, it is executed accordingly. However, in the absence of a matching control script, the LLM undertakes the planning process and generates a new control script to invoke AI modules and IoT devices.
The system performs preliminary feasibility validation on the newly generated script, checking for syntax errors or hardware-related issues. If the script fails to execute correctly, the system updates the context (coordinator’s short-term memory), gathering error information and triggering the LLM to regenerate the script based on the updated context. If the script continues to fail after multiple retries (more than three attempts), the system collects error information and reports it to the user. Conversely, if the script executes successfully, upon completion, the system summarizes the intermediate results gathered during script execution and generates an appropriate response based on the user’s instructions.
II-A User Interface
@@
II-B LLM
Within our framework, the LLM performs two key functionalities. First, it engages in conversations with users, generating contextually appropriate responses. Second, it does the planning and generates control scripts to address the tasks raised by the user, utilizing specialized AI modules and IoT devices as part of the available resources and methods. Python was selected as the programming language for our demo system.
II-B1 Conversation
To enhance the lifelike and immersive response of the LLM, we adopt a role-playing technique in which the LLM takes on the persona of a housekeeper. This approach is supported by carefully crafted system-level prompts, shown as follows. It is important to note that in order to provide an optimal user experience, specific user-related details may be required.
@@ II-B1 Conversation
II-B2 Script Generation
The LLM itself does not directly interact with AI modules and IoT devices. Instead, it generates Python control scripts to invoke AI modules and send control commands to IoT devices through network connections. Within this framework, we have proposed a Language-Code transformation scheme based on a finite-state machine (FSM) for generating scripts. Our Language-Code transformation scheme tackles two critical issues.
@@ II-B2 Script Generation
To address the first issue, the scheme enriches the information in the human command by leveraging knowledge about the environment. For instance, when a user instructs the agent to fetch a bottle of iced Coke, the agent should understand that iced Cokes are stored in the refrigerator, and the refrigerator is located in the kitchen corner, even if not explicitly mentioned. In our system, we enrich the command’s information with pre-stored environment knowledge using in-context learning.
Second, the LLM must precisely transform the enriched information (expressed in natural language) into executable code. However, the language description can become lengthy after enrichment to enhance clarity, and as a result, the generated code may become excessively complex, posing a challenge to error-free execution. Directly using the LLM to generate complex code from long language descriptions is unreliable due to its limited ability to handle intricate code with advanced functionality, as observed in [5]. This may result in the loss of crucial information from the human command or render the generated code unusable.
@@
Compared to the “Language-Code” transformation, our novel “Language-FSM-Code” transformation scheme maintains the same level of precision in conveying information since an FSM can precisely and comprehensively describe a task [6]. Additionally, by providing APIs that enable the LLM to reference and interact with AI modules and IoT devices, the complexity of sub-task transformation within each FSM state is manageable for current LLMs. Notably, our scheme enhances reliability without causing information loss.
@@
Fig. 4 depicts the Language-Code transformation procedure, which consists of three phases: First, natural language enrichment; next, Language-FSM transformation; and finally FSM-Code transformation. Our scheme utilizes pre-stored knowledge, including environment information, API documentation of AI modules and IoT devices, and example codes, to aid in the transformation.
@@
In phase one, the system enriches the task description using the environment information stored in the short-term memory of the coordinator. This information encompasses a list of available AI modules and IoT devices, their functional descriptions, as well as IoT device locations. The environment information enables a precise and executable task description.
@@
Phase two leverages the enriched task description and API information to generate the FSM representation of the task. The provision of API information helps the LLM to gain a better understanding of the capabilities of AI modules and IoT devices, thereby facilitating the development of a reasonable and efficient FSM design. Without API information, the FSM representation may include overly complex sub-tasks or redundant processes, causing inefficiencies.
@@
In phase three, the FSM-Code transformation is divided into three sub-phases. An FSM has two components: state operation and state transition. We generate code descriptions for each component in separate sub-phases, merging them in the final sub-phase. Our method simplifies the LLM’s task and improves reliability compared to a single-round interaction.
@@
In the first sub-phase, we provide the LLM with the FSM derived in phase two and API information. This enables the LLM to generate state operation code that incorporates the desired functionality. In the second sub-phase, example codes are provided to assist the LLM in understanding Python-based state transitions. Finally, the LLM is provided with the FSM, code generated in the previous sub-phases, and supplementary reference codes to facilitate the production of the final code.
II-C Coordinator
In our framework, the coordinator serves as the central hub, facilitating the interconnection and seamless operation of various system components. With its pivotal role in managing the flow of information and tasks, the coordinator ensures the smooth functioning of complex AI systems. It comprises three essential parts: short-term memory, long-term memory, and script executor.
The short-term memory contains the contextual information provided to the LLM for both conversation and code-generation purposes. This information serves as a crucial resource for the AI system, providing the necessary knowledge and understanding of user interactions, system states, and task requirements. It comprises four main components:
  1. 1.
    Environment Information: The component includes a list of available AI modules and IoT devices with their function descriptions, as well as the installation locations of IoT devices. It enables the LLM to be aware of the available AI modules, IoT devices, and their capabilities.
  2. 2.
    API Descriptions: This component offers a comprehensive description of the API functions used to invoke AI modules and control IoT devices within the script, encompassing details such as inputs, outputs, specifications, expected time costs, and other relevant information. These details enable the LLM to effectively leverage the available functionality.
  3. 3.
    Chat History: It retains the record of previous conversations between the system and the user, allowing the LLM to refer back to previous interactions.
  4. 4.
    Execution Results and Error Reports: The short-term memory stores the execution results of generated scripts and error reports returned by the script executor. This information helps the LLM assess the outcome of script execution and identify issues or errors that may have occurred.
The context is updated in real-time based on user interactions and the execution status of control scripts, ensuring that it is readily available for intelligent decision-making and adaptive behavior within the AI system.

### v3

# LLMind: Orchestrating AI and IoT with LLM for Complex Task Execution

@@ Abstract
The exploration of large language models (LLMs) for task planning and IoT automation has recently gained significant attention. However, existing works suffer from limitations in terms of resource accessibility, complex task planning, and efficiency. In this paper, we present LLMind, an LLM-based AI agent framework that enables effective collaboration among IoT devices for executing complex tasks. Inspired by the functional specialization theory of the brain, our framework integrates an LLM with domain-specific AI modules, enhancing its capabilities. Complex tasks, which may involve collaborations of multiple domain-specific AI modules and IoT devices, are executed through a control script generated by the LLM using a Language-Code transformation approach, which first converts language descriptions to an intermediate finite-state machine (FSM) before final precise transformation to code. Furthermore, the framework incorporates a novel experience accumulation mechanism to enhance response speed and effectiveness, allowing the framework to evolve and become progressively sophisticated through continuing user and machine interactions.
Introduction
@@
Large Language Model (LLM) has recently attracted tremendous attention from both academia and industry. In addition to excellent language skills, LLM also exhibits a high level of reasoning and planning capability [1, 2, 3, 4, 5]. Moreover, the remarkable feature of in-context learning substantially broadens the range of potential application scenarios for LLM, sparking the design of diverse general-purpose AI agents [6, 7, 8, 9, 10, 11, 12, 13].
In particular, LLM presents new transformative opportunities for the IoT era. In this work, we propose LLMind, an AI agent framework that transforms the usage of conventional IoT devices by integrating LLM as an orchestrator for accomplishing complex tasks using them.
@@
Compared with scripted intelligence and traditional IoT device control methods, the LLM orchestrator brings unparalleled flexibility and empowers IoT devices with advanced intelligence, enabling the coordination of diverse IoT devices to collaboratively accomplish intricate tasks. By harnessing the in-context learning capability of the LLM, the agent can achieve swift adaptation and optimal resource utilization, leading to efficient task solutions [14, 15]. Consequently, users can effortlessly express their needs without concerns about operational details.
@@
Integrating LLM into IoT device control presents a number of challenges before its promising potentials can be realized. First, a general LLM cannot handle specialized tasks like object detection or facial recognition, while there are already many AI models excelling in various specific domains. Thus, a challenge lies in combining the power of specialized but fragmented AI modules with LLM to handle various complex tasks. Furthermore, due to LLM’s slow inference speed and high computational costs, it is crucial to investigate methods that can enhance the system’s response speed and overall efficiency. Another significant challenge involves the effective control of IoT devices by the LLM-based AI agent. The LLM must accurately learn the diverse functionalities and operational characteristics of various IoT devices. Additionally, the design of efficient interaction mechanisms that enable users to naturally and intuitively express their needs to the LLM-based AI agent presents a further challenge. Our framework effectively overcomes the aforementioned challenges.
@@
First, taking inspiration from the functional specialization theory of the brain, our framework integrates domain-specific AI modules with the LLM, which serves as the task allocation governor. This synthesis enables the general-purpose LLM to effectively execute specific subtasks by invoking specialized AI modules to accomplish an overall complex task, enhancing system flexibility and practicality. Moreover, by incorporating new AI modules, the AI agent acquires new capabilities and continues to evolve to become even more capable.
@@
Second, to enhance the response speed and overall efficiency of the agent, we utilize a semantic analysis NLP model to classify and locally store LLM-generated plans (control scripts) for specific user commands. By identifying similar instructions in historical records, even with varying tones or expressions, the agent can retrieve validated solutions, thus avoiding complete replanning. This approach enables the system to accumulate experience and progressively enhance its proficiency through interactions.
@@
Third, the agent leverages the LLM for complex task planning and the generation of the corresponding control script using a novel finite-state machine approach, enabling efficient and precise invocation of AI modules and IoT devices. This methodology significantly improves accuracy and success rates.
@@
We refer to the intelligence level attained by our framework as Artificial Complex Intelligence (ACI), which serves as a bridge between Artificial Narrow Intelligence (ANI) [16, 17] and Artificial General Intelligence (AGI) [18, 19, 17], as illustrated in Fig. 1. ACI integrates the capabilities of various ANI modules and undergoes continuous evolution towards AGI. Our framework goes beyond mere device control, providing a comprehensive blueprint for a future intelligent, integrated, and collaborative IoT device ecosystem empowered by LLM.
Overview of the framework design
@@
Our framework, depicted in Fig. 2, consists of five components: user interface, LLM, coordinator, AI modules, and IoT devices.
The user can communicate with the AI agent via social media platforms to provide instructions and assign tasks. Once the task is comprehended, the LLM generates a control script to orchestrate IoT devices and AI modules, facilitating the completion of the assigned task.
If an error occurs during script execution, the error message is used as feedback and temporarily stored as context for subsequent trials. The successfully executed script will be locally stored and associated with the corresponding user instructions, providing a valuable learned experience for future executions.
Upon successful execution of the script or in the case of failure after multiple attempts, the agent collects the results and generates a report for the user.
User interface of the framework
@@
Functionalities of the LLM within the framework
The LLM in our framework performs two key functionalities.
First, it engages in conversations with users, generating contextually appropriate responses. To enhance the lifelike and immersive response of the LLM, we adopt a role-playing technique based on the LLM’s in-context learning ability. In our demo system, the LLM assumes the persona of a housekeeper, seamlessly blending into our everyday scenarios.
Second, the LLM does the planning and generates control scripts to address the tasks raised by the user, utilizing specialized AI modules and IoT devices as part of the available resources and methods. Notably, the LLM itself does not directly interact with AI modules and IoT devices. Instead, it generates control scripts to invoke AI modules and send control commands to IoT devices through network connections. This design enables the system to adapt to different AI modules and IoT devices without requiring extensive modifications to the LLM itself, enhancing scalability and interoperability. Python was selected as the programming language for our demo system due to its wide adoption in the field of artificial intelligence.
Within this framework, we propose a Language-Code transformation scheme using a finite-state machine (FSM) as a strategic and justified step for generating scripts. The complexity of user tasks often necessitates breaking them down into manageable subtasks. The FSM acts as an intermediate step, providing a structured approach to connect and sequence these subtasks. This choice is inspired by the inherent ability of FSMs to represent complex processes through a set of states and transitions, ensuring a systematic and organized generation of control scripts.
Functionalities and composition of the coordinator
The coordinator in our framework serves as the central hub, facilitating the interconnection and seamless operation of various system components. Without a central coordinator, the system components would operate in isolation, leading to inefficiencies and potential conflicts. The coordinator ensures smooth functioning of the agent system. It comprises three major parts: context repository, experience archive, and script executor.
The context repository contains the contextual information provided to the LLM for both conversation and code-generation purposes. This information serves as a crucial resource for the AI system, providing the necessary knowledge and understanding of user interactions, system states, and task requirements. It comprises four main components:
  1. 1.
    Environment Information: A list of available AI modules and IoT devices with their function descriptions, as well as the installation locations of IoT devices. This enables the LLM to be aware of the available AI modules, IoT devices, and their capabilities.
  2. 2.
    API Descriptions: A comprehensive description of the API functions used to invoke AI modules and control IoT devices within the script, encompassing details such as inputs, outputs, specifications, expected time costs, and other relevant information. This enables the LLM to effectively leverage the available functionality.
  3. 3.
    Chat History: Records of previous conversations between the system and the user, allowing the LLM to refer back to previous interactions.
  4. 4.
    Execution Results and Error Reports: Execution results of generated scripts and error reports returned by the script executor. This information helps the LLM assess the outcome of script execution and identify issues or errors that may have occurred.
The context is updated in real-time based on user interactions and the execution status of control scripts, ensuring that it is readily available for intelligent decision-making and adaptive behavior within the AI agent system.
The experience archive serves as a database for previously generated scripts. Valid scripts are stored locally and linked to their corresponding user instructions. This enables the system to reuse preexisting scripts when a similar instruction is given by the user in the future, resulting in improved efficiency and faster response speed. Importantly, the agent’s ability continues to grow through the accumulation and generalization of experience.
The control script generated by the LLM is passed to the executor, which provides the necessary runtime environment and offers API functions that can be called within the scripts. Furthermore, the executor monitors hardware status, handles runtime data, generates execution reports, and updates both the context repository and the experience archive.
The significance of introducing AI modules
Currently, LLM cannot effectively handle many domain-specific tasks [20]. To address this limitation, we integrated domain-specific AI modules into the framework, aiming to enhance overall performance by assigning specialized tasks to dedicated modules, resulting in improved capabilities, accuracy, and performance beyond LLM’s general knowledge.
The individual AI modules can be well versed in various specialized AI techniques, such as computer vision, natural language processing, speech recognition, sentiment analysis, recommendation systems, and more [21, 22, 23, 24, 25, 26, 27, 28, 29]. Each module is trained and optimized for its specific task, enabling it to provide highly accurate and specialized results.
Overall, the aggregate of the LLM and specialized AI modules, orchestrated by a coordinator, forms a general-purpose AI system capable of performing complex tasks. In a way, the system mimics the fact that the biological brain is also divided into coordinated parts with specialized functions.
Enabling efficient and secure control of IoT devices
IoT devices encompass a wide range of physical objects embedded with sensors, actuators, and connectivity features that enable internet connectivity and data exchange [30, 31]. Examples include smart appliances, wearables, industrial equipment, and environmental sensors.
To ensure efficient control of diverse IoT devices in our proposed framework, manufacturers must expose a set of well-documented API functions. These functions should encapsulate the necessary functionality, enabling the AI agent to initiate specific actions or retrieve information from the IoT devices.
It is important to emphasize that manufacturers should prioritize security considerations when exposing API functions for remote control [32]. By implementing robust authentication mechanisms, access controls, and encryption protocols [33, 34, 35, 36] within the API functions, they can ensure that only authorized entities have control over the devices. This approach guarantees secure and protected communications, safeguarding against unauthorized access or tampering.
Experiment configurations
We deployed the coordinator and AI modules of the system on an edge server. The edge server is connected to the same local network as the IoT devices via WiFi and can interact with ChatGPT using OpenAI’s APIs [37].
The available IoT devices in the system are as follows:
  1. 1.
  2. 2.
    TurtleBot mobile robot [38]: The TurtleBot mobile robot can receive target position coordinates through a network connection and autonomously plan paths to navigate within the room, intelligently avoiding obstacles.
  3. 3.
Additionally, the system integrates the following AI modules:
  1. 1.
  2. 2.
Check-in and security scenario experiment
Guests arrive at a party. The user sends a message to the AI agent: “Can you help me count the number of people in the room and identify who they are?” [^1] Demo video 1: https://youtu.be/3Al5qRntEEU Table I provides the dialogue content between different roles in the experiment. Among them, the LLM acts as the housekeeper, while the coordinator serves as the assistant.
Num. Conversations Contents
Please give me a moment to gather the information.
3 Assistant $\rightarrow$ Housekeeper Do you require any assistance?
In the absence of a valid script in the experience archive, the LLM generates a script, which is then executed in the executor.
There are also 2 people that can not be identified.
TABLE I: Dialog contents between different roles
In step 3, the coordinator questions the LLM about the need for assistance. Following each round of prompt and response interaction between the user and the LLM (similar to steps 1 and 2), the coordinator repeats the same question to the LLM. In the scenario where the LLM assumes the role of a remote housekeeper who relies on assistants in the user’s residence to carry out tasks. This inquiry serves to verify if the user has given specific instructions. Ambiguous instructions, as well as unrelated casual conversations, will cause the system to prompt the user to provide additional information or be filtered out. When the LLM indicates a need for help, it signifies that the user has assigned a task that could potentially be handled using AI modules and IoT devices. We employ a role-playing methodology that allows for the implementation of various roles and role-playing settings to facilitate the context-filtering process.
In steps 5 and 6, the coordinator prompts the LLM to generate a summary of the user’s instructions to facilitate the search for relevant historical scripts stored in the experience archive. This search process involves comparing the semantic content of the user’s instructions. However, the original task description may be poorly specified and be mixed with various linguistic habits. Additionally, the LLM may require multiple interactions with the user to fully grasp the details of specific tasks. To handle these conditions, in step 5, we request the LLM to provide a concise summary of the user’s instructions. Subsequently, we conduct a search for relevant scripts based on the LLM’s summary to enhance accuracy.
In this experiment, the coordinator could not find a suitable control script in the experience archive for the task. Consequently, the LLM is instructed to generate a new script. The code is generated using an FSM approach, as depicted in Fig. 3a. In addition to relying on the security camera installed on the ceiling, if the person cannot be identified due to low photo resolution or incomplete facial information, the AI agent will then instruct the TurtleBot to move to the position of the unknown person to take a close-shot photo for further assurance. The code generates two lists that store the execution results, one for known persons and another for unknown persons.
In step 7, once the script execution is complete, the LLM receives the execution results of the API functions. Throughout the script execution process, API functions may generate log information, encompassing intermediate results, warnings, and errors. This log information is collected and sent to the LLM as context for analysis.
(a) Check-in and security scenario
(b) Network management scenario
Figure 3: FSM of generated code
Network management scenario experiment
Our framework is designed to handle both one-off complex tasks and scenarios that involve multiple interactions with users. Next, we consider a scenario where a family of three individuals shares the same home WiFi network. [^2] Demo video 2: https://youtu.be/aTGD8EjQ8kM The interaction process between the AI agent and the user in this scenario is similar to the one described in the check-in and security scenario.
It is important to note that in this experimental scenario, we limit the total bandwidth of the shared network to 100 MB/s for demonstration purposes. Additionally, the speed limit for each user is fixed in a stepped manner, categorized as Low Rate (20 MB/s), Normal (30 MB/s), and High Rate (50 MB/s). In the first attempt, the AI agent successfully uses the network management API to elevate the user’s bandwidth from the Low Rate to the Normal level without exceeding the total bandwidth of the shared network. It then informs the user about the adjustment result.
However, if Eason remains dissatisfied with the network speed, he can ask, “Can you increase my internet speed once more?” Since the script for “improving internet speed for Eason” is already stored in the experience archive, the previously generated script is executed directly. However, in this experiment, during the execution of the script, it is determined that further increasing the network speed for Eason is not possible. This limitation arises because another user, Ada, is already in the High Rate list. Increasing the network speed for Eason again would exceed the total bandwidth limit of the shared network. Therefore, this execution is terminated, and the reason is communicated to Eason. The second scenario demonstrates that our framework is capable of handling tasks that involve multiple interactions with users.
Conclusion
The integration of LLMs into IoT device control offers transformative possibilities for developing intelligent IoT ecosystems. Our framework seamlessly combines LLMs’ language skills, reasoning abilities, and in-context learning with specialized AI modules, enhancing the flexibility and practicality of IoT systems. The proposed approaches, including FSM-based control script generation, user interaction, and historical script retrieval mechanisms, contribute to enhancing user experiences and system efficiency, with the system becoming increasingly intelligent as it learns from interactions among LLMs, humans, AI modules, and IoT devices. This comprehensive blueprint for intelligent and collaborative IoT ecosystems sets the stage for future advancements, where LLMs and AI-driven solutions enhance device coordination and convenience in our daily lives harmoniously.
Methods
Workflow
The operational workflow of a demo system is illustrated in the flowchart presented in Fig. 4.
Figure 4: System operational flowchart
Users can interact with the system by sending text messages through social media software. The LLM promptly acknowledges the receipt of the message. For casual chats and greetings, the system provides text replies without further processing. However, for specific requests and commands, the system formulates plans and strives to fulfill the user’s instructions using integrated AI modules and connected IoT devices.
After receiving a user instruction, the system strives for efficiency by first searching for validated and feasible historical scripts stored locally in the coordinator’s experience archive. This process utilizes NLP techniques to locate previously issued user commands with similar semantic meanings. If a matching control script is found, it is executed accordingly. However, in the absence of a matching control script, the LLM undertakes the planning process and generates a new control script to invoke AI modules and IoT devices.
The system performs preliminary feasibility validation on the newly generated script, checking for syntax errors or hardware-related issues. If the script fails to execute correctly, the system updates the context (the coordinator’s context repository), gathers error information, and triggers the LLM to regenerate the script based on the updated context. If the script continues to fail after multiple retries (more than three attempts), the system collects error information and reports it to the user. Conversely, if the script executes successfully, upon completion, the system summarizes the intermediate results gathered during script execution and generates an appropriate response based on the user’s instructions.
LLM role-playing
The LLM role-playing approach is supported by carefully crafted system-level prompts, as shown in Algorithm 1. It is important to note that in order to optimize the user experience, distinct user profiles are necessary for tailoring responses to individual users. These user-related details enable the system to provide personalized interactions, catering precisely to each user’s preferences and needs.
Algorithm 1 System-level prompts for role-playing
@@ Methods
FSM-based Language-Code transformation using LLM
Our Language-Code transformation scheme tackles two critical issues.
@@
To address the first issue, the scheme expands the information in the human command by leveraging knowledge about the environment. For instance, when a user instructs the agent to fetch a bottle of iced Coke, the agent should understand that iced Cokes are stored in the refrigerator, and the refrigerator is located in the kitchen corner, even if not explicitly mentioned. In our system, we expand the command’s information with pre-stored environment knowledge using in-context learning.
Second, the LLM must precisely transform the expanded information (expressed in natural language) into executable code. However, the language description can become lengthy after expansion to enhance clarity, and as a result, the generated code may become excessively complex, posing a challenge to error-free execution. Directly using the LLM to generate complex code from long language descriptions is unreliable due to its limited ability to handle intricate code with advanced functionality, as observed in [41]. This may result in the loss of crucial information from the human command or render the generated code unusable.
@@
Compared to the “Language-Code” transformation, our novel “Language-FSM-Code” transformation scheme maintains precision and conciseness in conveying information since an FSM can precisely and comprehensively describe a task [42, 43]. Additionally, by providing APIs that enable the LLM to reference and interact with AI modules and IoT devices, the complexity of sub-task transformation within each FSM state is manageable for current LLMs. Notably, our scheme enhances reliability without causing information loss.
@@
Fig. 5 depicts the Language-Code transformation procedure, which consists of three phases: First, context expansion; next, Language-FSM transformation; and finally FSM-Code transformation. Our scheme utilizes pre-stored knowledge, including environment information, API documentation of AI modules and IoT devices, and example codes, to aid in the transformation.
@@
For context expansion, the system expands the task description using the environment information stored in the context repository of the coordinator. This information encompasses a list of available AI modules and IoT devices, their functional descriptions, as well as IoT device locations. The environment information enables a precise and executable task description.
@@
The Language-FSM transformation leverages the expanded task description and API information to generate the FSM representation of the task. The provision of API information helps the LLM to gain a better understanding of the capabilities of AI modules and IoT devices, thereby facilitating the development of a reasonable and efficient FSM design. Without API information, the FSM representation may include overly complex sub-tasks or redundant processes, causing inefficiencies.
@@
The FSM-Code transformation is divided into three sub-phases. An FSM has two components: state operation and state transition. We generate code descriptions for each component in separate sub-phases, merging them in the final sub-phase. Our method simplifies the LLM’s task and improves reliability compared to a direct Language-Code transformation.
@@
In the first sub-phase, we provide the LLM with the FSM derived in the Language-FSM transformation and API information. This enables the LLM to generate state operation code that incorporates the desired functionality. In the second sub-phase, example codes are provided to assist the LLM in understanding Python-based state transitions. Finally, the LLM is provided with the FSM, code generated in the previous sub-phases, and supplementary reference codes to facilitate the production of the final code.
Historical script retrieval mechanism